# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 1.3 Ansatz
Ansatz 1.8 (Target).
Our target is to prove that the predicate
$\mathsf{CubicEqSat}(e):\equiv\exists\vec{x}\;P_{e}(\vec{x})=0,$
where $e$ codes a single polynomial of total degree $\leq 3$ with its variable count as part of the code, is $\Sigma^{0}_{1}$-complete. Equivalently, satisfiability of single cubic equations over $\mathbb{N}$ with instance-dependent arity is undecidable. The argument proceeds in three conceptual layers:
@@
    Compilation. We construct a uniform primitive recursive compiler
    $\varphi\longmapsto\mathcal{S}_{\varphi}$ (1.28)
@@
    $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\;\Longleftrightarrow\;\mathsf{CubicSat}(\ulcorner\mathcal{S}_{\varphi}\urcorner),$ (1.29)
    where $\mathsf{CubicSat}$ denotes satisfiability of a single degree-$\leq 3$ equation coinciding with “theoremhood” inside a Hilbert–Ackermann System.
@@
    Constraining. The degree budget is controlled locally: finite proof traces are arithmetized via Carryless Pairing in the Fibonacci basis (Section 3), syntactic and inference checks compile to quadratic constraints (Section 4), and the only cubic terms arise when a linear selector activates a quadratic obligation. These remain separate through the local degree certification; only at the final step are they collapsed to the single cubic equation by two-channel aggregation (Section 4.8).
@@
    Metatheory. By the classic Diagonal argument, any total correct decider for $\mathsf{CubicSat}$ would compose with the compiler to decide $\mathsf{Thm}$, (Section 5) leading to contradiction. If $\mathsf{RA}$ could internally certify correctness of such a decider on the compiler image, the provability bridge would force $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, contradicting Gödel’s Second Incompleteness Theorem (Gödel’s 2nd henceforth) [Gödel, 1931; Kleene, 1952; Feferman, 1960; Boolos et al., 2012] (Section 5.3).
Ansatz 1.9 (Metalogic).
The development harmonizes two formal methodologies: $\mathsf{RA}$ as our theory, while proof ereasure inside Calculus of Inductive Constructions of Coquand and Huet [1988] acts as a second metalogical verification tool. Layer $1$ and $2$ of the argument are fully constructive: all maps are primitive recursive, equivalences are witnessed by explicit transformations, and the entire pipeline is mechanized in Rocq without axioms beyond weak aritmetics. The construction remains constructive as far as it can: classical reasoning enters only at Layer $3$, where the Diagonal Lemma produces a fixed-point sentence whose existence relies on the final discharge via Gödel’s 2nd. This boundary is not a limitation but a structural feature: the undecidability of cubic satisfiability is inherently in need of a classical impossibility result, and the point at which constructivity is exhausted marks precisely the logical depth of the claim. Interestingly, the pipeline does not appeal to MRDP [Matiyasevich, 1993]: the Diophantine representation of proof evaluation is constructed directly, with explicit degree control at every step, rather than inherited from the historic theorem.
Proposition 1.10 (Numeralwise representability).
@@
$f(\vec{n})=m$ $\;\Longrightarrow\;\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m}),$ (1.30)
$\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m})\wedge\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m^{\prime}})$ $\;\Longrightarrow\;m=m^{\prime}.$ (1.31)
@@
Proposition 1.11 (Instancewise checker–constraint equivalence).
@@
$\mathrm{check}(\bar{p},\bar{t})\;\Rightarrow\;\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x}),$ (1.32)
@@
$\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x})\;\Rightarrow\;\mathrm{check}(\bar{p},\bar{t})\text{ holds in }\mathbb{N}.$ (1.33)
In particular, for each standard proof code $p$ of $\varphi$, $\mathsf{RA}$ proves the existence of a satisfying assignment for the emitted constraint system.
By construction. Each initial function (zero, successor, projection) is definable in $\mathcal{L}_{\mathsf{RA}}$. Closure under composition and primitive recursion follows by induction on the defining scheme, using $\mathrm{I}\Delta_{0}$ for the recursive step and $\mathrm{B}\Sigma_{1}$ to bound intermediate values. We use these facts only for numeralwise representability: for each concrete tuple of numerals $\overline{\vec{n}}$, $\mathsf{RA}$ proves the corresponding instance of $\varphi_{f}$. We do not assert that $\mathsf{RA}$ proves $\forall\vec{x}\,\exists!y\,\varphi_{f}(\vec{x},y)$ for arbitrary primitive recursive $f$. ∎
@@ 2 Constructive Environment
2.1 Mechanization
Ansatz 2.1.
This section establishes the computational infrastructure, coding of syntax, the proof predicate with derivability conditions, and some prerequisites of our fixed point. The pipeline described in the Ansatz 1.9 is implemented in Gallina, the specification language of
By Curry–Howard [Sørensen and Urzyczyn, 2006], the implicational fragment with $\bot$ admits a “simply-typed” $\lambda$-definable interpretation: each axiom (K, S, EFQ) realizes as a closed term and Modus ponens corresponds to application. The development formalizes the Hilbert-style proof evaluator in $\mathtt{Kernel\_API}$, $\mathtt{R13}$ (Section 4.1), compiles proof traces to cubic Diophantine constraints (Section 4), and extracts a certified boolean kernel,
$\mathtt{cubic\_accepts}:\mathbb{N}\longrightarrow\mathbb{B},$ (2.34)
@@ 2.2 Repository Structure and Trust Base
    Verification and constraint compilation ($\mathtt{R04}$$\mathtt{R11}$). Hilbert formula syntax and the executable proof checker ($\mathtt{R04}$$\mathtt{R05}$), formula and sequence coding via Carryless Pairing ($\mathtt{R06}$), and constraint assembly ($\mathtt{R08}$$\mathtt{R11}$) mapping proof-checking conditions to polynomial equations. Axiom and target constraints have degree $\leq 2$; cubic terms arise only when degree-1 selector variables activate quadratic obligations, keeping every emitted constraint at degree $\leq 3$. The assembly theorem $\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ ($\mathtt{R11}$) certifies the uniform bound $\delta\leq 3$.
@@
    Reduction endpoint and proof theory ($\mathtt{R14}$$\mathtt{R19}$ and $\mathtt{P00}$$\mathtt{P06}$). The reduction core and code bridge ($\mathtt{R14}$$\mathtt{R15}$), the reduction nucleus $\mathtt{compiler\_primitive}$ with coded evaluator semantics ($\mathtt{R16}$), OCaml extraction ($\mathtt{R17}$), and stable API boundaries ($\mathtt{R18}$$\mathtt{R19}$). The proof-theory layer ($\mathtt{P00}$$\mathtt{P06}$) mechanizes concrete provability ($\mathtt{P00}$), HBL derivability conditions ($\mathtt{P01}$), the fixed-point construction ($\mathtt{P02}$), the premise-free coded undecidability theorem ($\mathtt{P03}$), the RA-certification corollary ($\mathtt{P04}$), and the toggle-witness contradiction core ($\mathtt{P05}$$\mathtt{P06}$).
Ansatz 2.3.
@@
    $\mathtt{cubic\_accepts}:\mathsf{Proof}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\},$ (2.35)
    the proof-checking kernel. Internally the module re-derives its own bool, nat, list, and form types and contains the full verification pipeline: structural formula equality (form_eqb), axiom-schema recognizers (is_K, is_S, is_EFQ), bounded Modus ponens witness search (mp_witness), and line-by-line validation (check_lines). The entry point $\mathtt{cubic\_accepts}$ is definitionally equal to $\mathtt{check}$; the kernel API ($\mathtt{R13}$) re-exports the checker under a reduction-facing alias.
@@
    cubic_compile.ml exports a (legacy) polynomial,
    $\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial})$ (2.36)
@@
    $\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{Polynomial}\times\mathsf{Polynomial},$ (2.37)
@@
    $\mathtt{Expr}(\mathtt{Const}\mid\mathtt{Var}\mid\mathtt{Add}\mid\mathtt{Sub}\mid\mathtt{Mult})$ (2.38)
@@
    $\mathtt{degree}:\mathtt{Expr}\to\mathbb{N}$ (2.39)
    Given a proof and target, $\mathtt{emit\_cubic\_system}$ produces polynomial constraints, while the the endpoint $\mathtt{emit\_single\_cubic}$ applies two-channel aggregation and returns a single cubic equality surface $(P,Q)$ representing $P=Q$. Mapping $\mathtt{degree}\circ\mathtt{poly\_expr}$ over the emitted expressions empirically verifies the degree-$3$ bound of Proposition 4.28. The module additionally includes the constraint types ($\mathtt{C\_Axiom}$, $\mathtt{C\_MP}$, $\mathtt{C\_Justification}$, $\mathtt{C\_Target}$), the canonical assignment builder, and the Carryless Pairing and carryless sequence-indexing infrastructure required by the compilation pipeline.
@@
Intuition.
A clarification on the diagonal construction is necessary, since it appears at two distinct levels. The Rocq development contains a code-level diagonal ($\mathtt{P02}$$\mathtt{P03}$): it operates entirely on natural-number codes, producing a fixed-point code $g$ satisfying $g=\mathtt{apply\_template}\;T\;g$ by arithmetic on encodings, and derives $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$)—the impossibility of a coded decider—constructively, without classical axioms. This is a theorem about codes: it quantifies over natural numbers representing programs, not over Turing Machines as external objects. The classical meta-theoretic diagonal of Section 5.2 (Theorem 5.2) is a separate argument: it applies the Diagonal Lemma to arbitrary Turing Machines with case analysis via the Law of the Excluded Middle, and is not Rocq-verified. Likewise, Gödel’s 2nd as applied to $\mathsf{RA}$ (Corollary 5.7) relies on the arithmetization of Section 3.4 and standard classical logic, and is presented as an ordinary mathematical argument. A machine-readable cross-reference map of manuscript claims to Rocq theorems is provided in $\mathtt{TechnicalSummary.yaml}$.
Remark.
The reduction used for $\Sigma^{0}_{1}$-hardness does not fix a proof-length bound. The compiler $f(u)$ treats the entire proof witness as a single carryless list code $p$ (together with its per-line selectors), and emits one cubic equation whose free variables include $p$, the target code $u$, and all selector/auxiliary variables. The Diophantine instance is therefore
$\exists\,p,\,\vec{x}\;\Phi(u,p,\vec{x})=0$
@@
$\operatorname{Prov}_{\mathsf{RA}}(u)\;\Longleftrightarrow\;\exists\,p,\vec{x}\;\Phi(u,p,\vec{x})=0\;\Longleftrightarrow\;\mathsf{CubicEqSat}(f(u)),$
where $f$ is primitive recursive and $\mathsf{CubicEqSat}$ ranges over a single cubic equation with fixed variable budget.
The banded aggregation uses dynamically sized Fibonacci bands: the width parameter $K$ is computed from the size of the emitted constraint system (number of constraints / channels), not from a fixed global constant. Each channel is paired with a capacity constraint of the form $B_{K}-1-A_{i}=\sum_{j=1}^{4}s_{ij}^{2}$, where $B_{K}=F_{K+1}$ and $s_{ij}$ are slack variables. Thus any assignment with $A_{i}\geq B_{K}$ is rejected. Since every $A_{i},B_{i}$ is a polynomial in the proof/selector variables of degree $\leq 2$, the capacity bound is computable from the instance size, and the added slack variables preserve equivalence: a satisfying assignment to the original system extends to one satisfying the capacity equations, and any violation of the bound is ruled out by the slack equality.

### v7

@@ 1.3 Ansatz
Ansatz 1.3.1 (Target).
We fix the following as the canonical decision predicate throughout:
$\mathsf{CubicEqSat}(e):\equiv\exists\vec{x}\;P_{e}(\vec{x})=0,$ (28)
where $e$ codes a single polynomial of total degree $\leq 3$ with its variable count as part of the code. All completeness and undecidability statements in this paper are formulated for $\mathsf{CubicEqSat}$ in the $R=0$ form; alternative surfaces (such as the pair surface $P=Q$ used in the mechanized kernel) are connected by Lemma 2.2.3. Our target is to prove that $\mathsf{CubicEqSat}$ is $\Sigma^{0}_{1}$-complete. Equivalently, satisfiability of single cubic equations over $\mathbb{N}$ with instance-dependent arity is undecidable. The argument proceeds in three conceptual layers:
@@
    We construct a uniform primitive recursive compiler
    $\varphi\longmapsto\mathcal{S}_{\varphi}$ (29)
@@
    $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\;\Longleftrightarrow\;\mathsf{CubicSat}(\ulcorner\mathcal{S}_{\varphi}\urcorner),$ (30)
    where $\mathsf{CubicSat}$ denotes satisfiability of a single degree-$\leq 3$ equation coinciding with “theoremhood” inside a Hilbert–Ackermann System [Hilbert and Ackermann, 1938].
@@
    The degree budget is controlled locally: finite proof traces are arithmetized via Carryless Pairing in the Fibonacci basis (Section 3), syntactic and inference checks compile to quadratic constraints (Section 4), and the only cubic terms arise when a linear selector activates a quadratic obligation. These remain separate through the local degree certification; only at the final step are they collapsed to the single cubic equation by two-channel aggregation (Section 4.8).
@@
    By the classic Diagonal argument, any total correct decider for $\mathsf{CubicSat}$ would compose with the compiler to decide $\mathsf{Thm}$ (Section 5), leading to contradiction. If $\mathsf{RA}$ could internally certify correctness of such a decider on the compiler image, the provability bridge would imply
    $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA}),$ (31)
    contradicting Gödel’s Second Incompleteness Theorem (Gödel’s 2nd henceforth) [Gödel, 1931; Kleene, 1952; Feferman, 1960; Boolos et al., 2012] (Section 5.3).
Ansatz 1.3.2 (Automated type checking as metalogic).
The development uses $\mathsf{RA}$ as object theory and proof erasure specified by the Calculus of Inductive Constructions [Coquand and Huet, 1988] as metalogical verification. Layers $1$ and $2$ are fully constructive: all maps are primitive recursive, all equivalences explicitly witnessed, and the pipeline is mechanized in Rocq without axioms beyond weak arithmetics. Classical reasoning enters only at Layer $3$, where the Diagonal Lemma produces a fixed-point sentence discharged via Gödel’s 2nd. The pipeline does not appeal to MRDP [Matiyasevich, 1993]: the Diophantine representation of proof evaluation is constructed directly, with explicit degree control at every step.
Proposition 1.3.3 (Numeralwise representability).
@@
$f(\vec{n})=m\;\Longrightarrow\;\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m}),$ (32)
$\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m})\wedge\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m^{\prime}})\;\Longrightarrow\;m=m^{\prime}.$ (33)
@@
Proposition 1.3.4 (Instancewise checker–constraint equivalence).
@@
$\mathrm{check}(\bar{p},\bar{t})\;\Rightarrow\;\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x}),$ (34)
@@
$\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x})\;\Rightarrow\;\mathrm{check}(\bar{p},\bar{t})\text{ holds in }\mathbb{N}.$ (35)
In particular, for each standard proof code $p$ of $\varphi$, $\mathsf{RA}$ proves the existence of a satisfying assignment for the emitted constraint system (Table 1).
By construction. Each initial function (zero, successor, projection) is definable in $\mathcal{L}_{\mathsf{RA}}$. Closure under composition and primitive recursion follows by induction on the defining scheme, using $\mathrm{I}\Delta_{0}$ for the recursive step and $\mathrm{B}\Sigma_{1}$ to bound intermediate values. We use these facts only for numeralwise representability: for each concrete tuple of numerals $\overline{\vec{n}}$, $\mathsf{RA}$ proves the corresponding instance of $\varphi_{f}$. We do not assert that $\mathsf{RA}$ proves
$\forall\vec{x}\,\exists!y\,\varphi_{f}(\vec{x},y)$ (36)
for arbitrary primitive recursive $f$. ∎
@@ 2 Constructive Environment
2.1 Mechanization in Rocq
Mechanization 2.1.1 (Proof environment T002).
This section establishes the computational infrastructure, coding of syntax, the proof predicate with derivability conditions, and some prerequisites of our fixed point. The pipeline described in the Ansatz 1.3.2 is implemented in Gallina, the specification language of
By Curry–Howard Correspondence [Sørensen and Urzyczyn, 2006], the implicational fragment with $\bot$ admits a “simply-typed” $\lambda$-definable interpretation: each axiom (K, S, EFQ) realizes as a closed term and modus ponens corresponds to application. The development formalizes the Hilbert-style proof evaluator in $\mathtt{Kernel\_API}$, $\mathtt{R13}$ (Section 4.1), compiles proof traces to cubic Diophantine constraints (Section 4), and extracts a certified boolean kernel,
$\mathtt{cubic\_accepts}:\mathbb{N}\longrightarrow\mathbb{B},$ (37)
@@ 2.2 Repository Structure and Trust Base
    Verification and constraint compilation ($\mathtt{R04}$$\mathtt{R06}$, $\mathtt{R08}$$\mathtt{R11}$). Hilbert-formula syntax and the executable proof checker ($\mathtt{R04}$$\mathtt{R05}$), formula and sequence coding via Carryless Pairing ($\mathtt{R06}$), and constraint assembly ($\mathtt{R08}$$\mathtt{R11}$) mapping proof-checking conditions to polynomial equations. Axiom and target constraints have degree $\leq 2$; cubic terms arise only when degree-1 selector variables activate quadratic obligations, keeping every emitted constraint at degree $\leq 3$. The assembly theorem $\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ ($\mathtt{R11}$) certifies the uniform bound $\delta\leq 3$.
@@
    Reduction endpoint and proof theory ($\mathtt{R14}$$\mathtt{R19}$, with auxiliary $\mathtt{R20}$$\mathtt{R22}$, and $\mathtt{P00}$$\mathtt{P06}$). The reduction core and code bridge ($\mathtt{R14}$$\mathtt{R15}$), the reduction nucleus $\mathtt{compiler\_primitive}$ with coded evaluator semantics ($\mathtt{R16}$), OCaml extraction ($\mathtt{R17}$), and stable API boundaries ($\mathtt{R18}$$\mathtt{R19}$). Auxiliary modules cover aggregated evaluation and channel/pairing support ($\mathtt{R20}$$\mathtt{R22}$). The proof-theory layer ($\mathtt{P00}$$\mathtt{P06}$) mechanizes concrete provability ($\mathtt{P00}$), HBL derivability conditions ($\mathtt{P01}$), the fixed-point construction ($\mathtt{P02}$), the premise-free coded undecidability theorem ($\mathtt{P03}$), the RA-certification corollary ($\mathtt{P04}$), and the toggle-witness contradiction core ($\mathtt{P05}$$\mathtt{P06}$).
Mechanization 2.2.2 (Extraction).
@@
    $\mathtt{cubic\_accepts}:\mathsf{Proof}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\},$ (38)
    the proof-checking kernel. Internally the module re-derives its own bool, nat, list, and form types and contains the full verification pipeline: structural formula equality (form_eqb), logical axiom-schema recognizers (is_K, is_S, is_EFQ), non-logical $\mathsf{RA}$ axiom-schema recognizers (successor axioms, $\Delta_{0}$-induction instances, $\mathrm{B}\Sigma_{1}$-collection instances), bounded modus ponens witness search (mp_witness), and line-by-line validation (check_lines). The entry point $\mathtt{cubic\_accepts}$ is definitionally equal to $\mathtt{check}$; the kernel API ($\mathtt{R13}$) re-exports the checker under a reduction-facing alias.
@@
    cubic_compiler.ml exports a (legacy) polynomial,
    $\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial})$ (39)
@@
    $\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{Polynomial}\times\mathsf{Polynomial},$ (40)
@@
    $\mathtt{Expr}(\mathtt{Const}\mid\mathtt{Var}\mid\mathtt{Add}\mid\mathtt{Sub}\mid\mathtt{Mult})$ (41)
@@
    $\mathtt{degree}:\mathtt{Expr}\to\mathbb{N}$ (42)
    Given a proof and target, $\mathtt{emit\_cubic\_system}$ produces polynomial constraints, while the endpoint $\mathtt{emit\_single\_cubic}$ applies two-channel aggregation and returns a single cubic equality surface $(P,Q)$ representing $P=Q$. Mapping $\mathtt{degree}\circ\mathtt{poly\_expr}$ over the emitted expressions empirically verifies the degree-$3$ bound of Proposition 4.8.10. The module additionally includes the constraint types ($\mathtt{C\_Axiom}$, $\mathtt{C\_MP}$, $\mathtt{C\_Justification}$, $\mathtt{C\_Target}$), the canonical assignment builder, and the Carryless Pairing and sequence-indexing infrastructure required by the compilation pipeline.
    Surface convention. The kernel API exposes the pair surface $(P,Q)$ to avoid introducing subtraction into the internal AST. For the canonical Diophantine coding used in all completeness statements (Equation 28), the compiler output is collapsed to a single polynomial equation via PQ=0P-Q=0 (degree unchanged by Lemma 2.2.3). The two surfaces are definitionally intertranslatable; see Lemma 2.2.3.
@@
Lemma 2.2.3 (Interface equivalence).
There are primitive recursive, degree-preserving translations
$\mathsf{toZero}(P,Q)$ $:=P-Q,$ (43)
$\mathsf{toPair}(R)$ $:=(R,\,0),$ (44)
such that for all assignments $\rho$ over $\mathbb{N}$:
$P(\rho)=Q(\rho)\;\Longleftrightarrow\;\mathsf{toZero}(P,Q)(\rho)=0,$ (45)
$R(\rho)=0\;\Longleftrightarrow\;\mathsf{toPair}(R)(\rho)=(R(\rho),\,0).$ (46)
In particular, both encodings yield the same satisfiability predicate up to a primitive recursive many-one equivalence:
$\deg(P-Q)\leq\max(\deg P,\deg Q)$ (47)
whenever $P\neq Q$, and
$\deg R=\max(\deg R,0).$ (48)
The Rocq artifact pins the pair surface $(P,Q)$ because it is convenient for normalization and avoids subtraction in the internal AST; the proof-of-concept and all manuscript theorems use the $R=0$ surface (Equation 28).
Immediate. $P=Q$ iff PQ=0P-Q=0, and $R=0$ iff $(R,0)$ is a pair with equal components. Both translations are compositional, primitive recursive in the polynomial codes, and preserve total degree. ∎
Mechanization 2.2.4 (Fixed points).
The diagonal construction appears at two distinct levels:
  1. (i)
    Code-level ($\mathtt{P02}$$\mathtt{P03}$). Operates on natural-number codes, producing a fixed-point code $g$ with $g=\mathtt{apply\_template}\;T\;g$, and derives $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) constructively, without classical axioms.
  2. (ii)
    Meta-theoretic (Section 5.2, Theorem 5.2.1). Applies the Diagonal Lemma to arbitrary Turing Machines via the Law of the Excluded Middle; not Rocq-verified. Gödel’s 2nd as applied to $\mathsf{RA}$ (Corollary 5.3.5) likewise uses standard classical logic.
A cross-reference map of manuscript claims to Rocq theorems is provided in $\mathtt{TechnicalSummary.yaml}$.
Remark 2.2.5.
The reduction used for $\Sigma^{0}_{1}$-hardness does not fix a proof-length bound. The compiler $f(u)$ treats the entire proof witness as a single carryless-list code $p$ (together with its per-line selectors), and emits one cubic equation whose free variables include $p$, the target code $u$, and all selector/auxiliary variables. The Diophantine instance is therefore
$\exists\,p,\,\vec{x}\;\Phi(u,p,\vec{x})=0$ (49)
@@
$\operatorname{Prov}_{\mathsf{RA}}(u)\;\Longleftrightarrow\;\exists\,p,\vec{x}\;\Phi(u,p,\vec{x})=0\;\Longleftrightarrow\;\mathsf{CubicEqSat}(f(u)),$ (50)
where $f$ is primitive recursive and $\mathsf{CubicEqSat}$ ranges over a single cubic equation with fixed variable budget. The banded aggregation uses dynamically sized Fibonacci bands: the width parameter $K$ is computed from the size of the emitted constraint system (number of constraints / channels), not from a fixed global constant. Each channel is paired with a capacity constraint of the form
$B_{K}-1-A_{i}=\sum_{j=1}^{4}s_{ij}^{2},$ (51)
where
$B_{K}=F_{K+1}\quad\text{and}\quad s_{ij}$ (52)
are slack variables, hence any satisfying assignment automatically satisfies
$0\leq A_{i}<B_{K}\quad\text{and}\quad 0\leq B_{i}<B_{K},$ (53)
an assignment with $A_{i}\geq B_{K}$ is rejected because the slack equation has no nonnegative solution. Since every $A_{i},B_{i}$ is a polynomial in the proof/selector variables of degree $2$, the capacity bound is computable from the instance size, and the added slack variables preserve equivalence: a satisfying assignment to the original system extends to one satisfying the capacity equations, and any violation of the bound is ruled out by the slack equality. No external representation bound is needed; the base is determined syntactically and the digit bounds are enforced internally (Lemma 4.8.5).